# IberFire - a detailed creation of a spatio-temporal dataset for wildfire risk assessment in Spain

### v1

@@ 9 Cataluña 19 Melilla
All five geographic location information features remain invariant over time. Consequently, they were incorporated into the datacube only with spatial coordinates, without the time dimension.
@@ 2.5 Land usage: Corine Land Cover
The Copernicus Corine Land Cover (CLC) dataset [9] offers a standardized classification of land cover types across Europe, distinguishing 44 discrete categories. It represents the continent as a regular grid of 100m wide cells, assigning to each cell an integer value between 1 and 44 that corresponds to a specific land cover class.
@@ 44 Water bodies Marine waters Sea and ocean
The original spatial resolution of the CLC dataset is 100m $\times$ 100m; consequently, resampling of the data was needed to match the 1km $\times$ 1km resolution of the IberFire datacube. Using QGIS, the proportion of each of the 44 classes within each 1km $\times$ 1km cell was calculated. This resulted in 44 continuous features, denoted as CLC_i for $i=1,2,\dots,44$, with values ranging between 0 and 1.
Given that these features represent proportions and sum to 1 for each 1km $\times$ 1km cell, and considering the hierarchical structure of the CLC classification, five additional features were derived corresponding to the higher-level groupings defined in Label 1. Specifically, for each higher-level category, its proportion within a cell was computed by summing the relevant CLC_i variables that fall inside the category. For example, the proportion of Artificial surfaces was obtained by aggregating the values of CLC_1 through CLC_11.
@@
The CLC dataset is updated every six years; accordingly, the 2006, 2012, and 2018 editions were used in this work, as the 2024 version was not yet available at the time of writing. As previously discussed, time-invariant features in the datacube can be stored using only spatial dimensions. Given the low update frequency of the CLC dataset, including a temporal dimension would result in unnecessary duplication of values. Therefore, all CLC-derived features were stored without the time axis. To differentiate between editions, the corresponding year was appended to each feature name, as better described in Table 6 of Section 3. Each of the three CLC editions contributed 63 variables, this approach ultimately yielded $63\cdot 3=189$ distinct CLC-derived spatial features in the datacube.
@@ 2.6 Topography variables: European Digital Elevation Model
Topography is a well-established factor influencing the spread and intensity of forest fires. To capture these effects, topographic features were downloaded from the European Digital Elevation Model (EU-DEM) provided by OpenTopography [15]. The original EU-DEM dataset offers elevation values at a 30m×$\times$30m spatial resolution. Using this elevation data, additional variables such as slope, aspect and roughness of the terrain can be derived, all of which are also offered by OpenTopography.
However, the specific methods used by OpenTopography to calculate the slope and roughness are not documented, leaving the units of these variables undefined. In contrast, the aspect is expressed in degrees, ranging from 0°to 360°, indicating the direction in which the slope of each 30m×$\times$30m terrain cell faces, with 0° corresponding to the north and the values increasing clockwise.
The elevation, slope, roughness and aspect were downloaded and, for the first three features, the mean and standard deviation in each 1km×$\times$1km cell were calculated. Therefore, a total of 6 different features were obtained from elevation, slope, and roughness. For aspect, the feature was discretised into 8 different classes defined in Table 3. For each of the 1km×$\times$1km cell, the proportion of each class inside the cell was calculated, resulting in 8 new features called aspect_i for $i=1,\dots,8$. Additionally, an extra feature, aspect_NODATA, was included to represent the proportion of pixels within each 1km×$\times$1km cell for which no aspect value was available, which is highly correlated with is_waterbody. This lack of data typically occurs in areas covered by lakes and rivers, where aspect values are undefined due to the absence of terrain elevation gradients.
@@ 2.7 Human activity
It is widely recognised that most wildfires are directly related to human activities, either intentionally or accidentally [24]. To model this phenomenon, six human-related explanatory variables were considered: distance to roads, distance to waterways, distance to railways, designation within the Natura 2000 protected network, population density, and holiday periods. These variables were used to derive a total of 20 features, which were subsequently integrated into the datacube.
@@
The first two variables, distance to roads and distance to waterways, were obtained from WorldPop [16], which provides data at an approximate spatial resolution of 100m×$\times$100m, represented in kilometres. To upscale the data to the target resolution, the mean and standard deviation within each 1km×$\times$1km cell were computed using QGIS. These aggregated statistics were then incorporated into the datacube, resulting in four derived features from the original two variables. These features were considered invariant over time and therefore added with only spatial coordinates.
@@
For distance to railways, railway vectorial data in Spain were retrieved from OpenStreetMap [17]. Using QGIS, a 100m×$\times$100m raster layer was generated to represent the distance to the nearest railway geometry from each 100m×$\times$100m cell. Then, the same procedure was applied, and the average and standard deviation were calculated. These two features were then added to the datacube as spatial data.
@@
The Natura 2000 network, a European ecological network for biodiversity conservation, was included due to its relevance in fire data reported by EFFIS, which indicates the proportion of fire within these protected areas. Using vectorial boundaries of the Natura 2000 network [25], a binary raster layer was created with QGIS indicating whether each 1km×$\times$1km cell falls within the protected area. Since this variable remains constant over time, it was integrated as a spatial layer into the datacube.
@@
Population density was incorporated as a proxy for human presence and potential ignition sources. Annual data were obtained from WorldPop for the years 2008 to 2020 [16], due to the unavailability of more recent data. The original data were retrieved with a resolution of approximately 1km $\times$ 1km, and average resampling was applied to match the coordinate values of IberFire. Given the annual update frequency of the data, population density were stored as spatial features to prevent unnecessary duplication of values, mirroring the process conducted with the CLC-derived features. As a result, this variable contributed 13 distinct spatial layers to the final dataset.
@@
Finally, a binary feature representing holiday periods was included. This variable accounts for the increased presence of people in rural and natural areas during weekends and public holidays, potentially raising the risk of fire ignition. A cell was flagged as a holiday if the corresponding date was a Saturday, Sunday, or a national or regional public holiday in Spain. To determine public holidays, the Python library holidays [26] was utilized. This library provides holiday dates for various countries and their subdivisions, allowing for precise identification of holidays at the autonomous community level in Spain. Consequently, the AutonomousCommunities feature was employed to assign the appropriate regional holidays to each cell. The resulting is_holiday binary layer was added as a spatio-temporal feature within the datacube.
@@ 2.8 Meteorological variables: ERA5-Land
To account for these dynamics, data from ERA5-Land [10] were integrated, a high-resolution global reanalysis dataset provided by the Copernicus Climate Data Store. ERA5-Land offers hourly meteorological variables at a spatial resolution of 9 km $\times$ 9 km, from 1950 to the present. This dataset is particularly well-suited for environmental modelling due to its global coverage, temporal consistency, and availability of multiple curated atmospheric variables relevant to fire risk assessment. Within the Copernicus Climate Data Store, two data access modes are available: raw hourly observations and post-processed daily statistics. Both sources were used in the construction of the IberFire datacube, depending on the specific requirements of each variable.
Although ERA5-Land provides temporally consistent data, its availability is subject to a delay of five days, making it unsuitable for real-time fire risk prediction, where daily forecasts are required. Consequently, while ERA5-Land data were used to construct the datacube, real-time model deployment is intended to rely on meteorological station measurements. The use of measurements for the construction of the datacube was also considered, but the lack of historical data for many variables in various meteorological stations posed significant limitations.
@@
A total of 17 meteorological features were derived from ERA5-Land data. The features can be grouped as follows: temperature, relative humidity, surface pressure, precipitations, wind speed, and wind direction. This section outlines the methodology used to obtain and incorporate these variables into the IberFire datacube, including the selection of source variables and the transformations necessary to ensure compatibility with the format and units used in AEMET observations.
@@
Relative humidity is a critical variable for fire risk assessment, as it directly affects the moisture content of vegetation and, consequently, the likelihood of ignition and propagation. However, ERA5-Land does not provide this variable directly, while AEMET does. To address this, hourly relative humidity values were derived from two available ERA5-Land variables: the 2m temperature and the 2m dewpoint temperature. The last one represents the temperature at which the air becomes saturated with moisture, at 2 metres above the ground. Using these two temperature variables, relative humidity values were computed by applying the Magnus formula [27], an empirically validated approach for estimating saturation vapor pressure:
Relative $\textit{Relative Humidity}=\frac{\exp(\frac{17.625\cdot D_{p}}{243.04+D_{p}})}{\exp(\frac{17.625\cdot T}{243.04+T})}$ (1)
@@
From the computed hourly values, four daily statistics were derived and included in the datacube: the mean, minimum, maximum, and range of relative humidity. This process required downloading and processing the full hourly time series of both temperature-related variables. Notice that the relative humidity daily statistics could not be accurately calculated from the ERA5-Land temperature summary statistics. For example, the mean of relative humidity can not be calculated by introducing the mean of $D_{p}$ and $T$ on the equation (1). The same happens for the other statistics too.
@@
$\lVert\vec{u}+\vec{v}\rVert=\lVert\vec{u}\rVert+\lVert\vec{v}\rVert\,.$ (2)
@@
wind_speed_mean. The data were stored with the original ERA5-Land units, meters per second (m/s), since they match with the units used by AEMET.
@@
Wind direction is conventionally defined as the direction from which the wind originates, expressed in degrees measured clockwise from true north. AEMET provides two daily wind-direction metrics: the average wind direction and the direction at the daily maximum wind speed. Meanwhile, ERA5-Land offers hourly wind data at a resolution of 9km $\times$ 9km, represented by horizontal (u_wind) and vertical (v_wind) components. To convert these components into standard wind direction, equation (3) is employed, where $\alpha$ denotes the angle between $(u,v)$ and the north vector $(0,1)$.
$\textit{Wind direction}=(\alpha-180^{\circ})\bmod 360^{\circ}\,,\quad\alpha=\angle\bigl{(}(0,1),(u,v)\bigr{)}\,.$ (3)
@@ 2.9 Vegetation indices: Copernicus Land Monitoring Service
The last group of features integrated into the datacube is the vegetation indices, which are highly used for making fire-risk predictions, as they can represent the dryness of the plant life. Five different indices were added: FAPAR (Fraction of Photosynthetically Active Radiation), LAI (Leaf Area Index), NDVI (Normalized Difference Water Index), LST (Land Surface Index), and SWI (Soil Water Index). All of them were retrieved from the Copernicus Land Monitoring Service (CLMS) [11] at various spatial and temporal resolutions.
For each vegetation index, multiple data sources from the Copernicus Land Monitoring Service (CLMS) were employed, each differing in spatial and temporal resolution. Depending on the source, some datasets provide global coverage, while others are limited to the European continent. To ensure efficiency and relevance, the data were downloaded specifically for the region of interest illustrated in Figure 1. The following sections offer a detailed description of the sources selected for each index, the procedures followed during data acquisition, and the preprocessing steps applied to harmonise all variables with the target spatial resolution of 1km $\times$ 1km and daily temporal granularity required by the IberFire datacube. This process was done using Python, and all the vegetation indices were incorporated into the dataset as spatio-temporal features.
@@
The Fraction of Absorbed Photosynthetically Active Radiation (FAPAR) is a key biophysical variable that is commonly used as an indicator of vegetation health. It has been identified as one of the 50 Essential Climate Variables (ECVs) by the Global Climate Observing System (GCOS) [28]. In the context of wildfire risk prediction, FAPAR is particularly relevant as it reflects the photosynthetic activity and water stress level of vegetation, which are tightly linked to fuel dryness.
The CLMS provides multiple datasets for the FAPAR variable; however, no single source spans the entire temporal range required by the IberFire datacube. To achieve full temporal coverage, two complementary datasets were integrated. The first source offers FAPAR measurements at a spatial resolution of 1km $\times$ 1km with a 10-day temporal frequency and was used to cover the period from 01/12/2007 to 30/04/2020. The second source, used for the remaining period from 01/05/2020 to 31/12/2024, provides enhanced spatial resolution at 300m $\times$ 300m, while maintaining the same temporal frequency.
The two datasets were harmonized as follows. For the first source, since its native resolution matched the target resolution of the datacube, values were directly aligned to the grid using nearest-neighbour resampling. In contrast, the higher-resolution values from the second source were aggregated by computing the mean FAPAR value within each corresponding 1km $\times$ 1km cell, effectively applying the average resampling method.
@@
The download and preprocessing procedures for this variable followed the same approach as for the FAPAR index. A dataset with a 1 km spatial resolution was used for the period from 01/12/2007 to 30/04/2020 and was aligned to the datacube using nearest-neighbour resampling. For the remaining period, from 01/05/2020 to 31/12/2024, a dataset at 300 m resolution was employed, and values were aggregated to 1 km resolution using average resampling.
@@
The Normalized Difference Vegetation Index (NDVI) is a remote sensing indicator that estimates vegetation health. It is calculated from the red and near-infrared spectral bands and typically ranges from $-1$ to $1$, with higher values indicating denser and healthier vegetation. In the context of wildfire risk assessment, NDVI serves as a proxy for vegetation condition and fuel availability.
As with the previous indices, two complementary datasets from the CLMS were used to ensure full temporal coverage. The first dataset, with a spatial resolution of $1km\times 1km$ and a 10-day frequency, covers the period from 01/12/2007 to 30/06/2020. The second dataset, with a finer resolution of 300 m and the same temporal frequency, was used from 01/07/2020 to 31/12/2024. The harmonization process mirrored that used for FAPAR and LAI: nearest-neighbour resampling was applied to the 1 km dataset, while the 300 m data were aggregated to 1 km using average resampling.
@@
The Land Surface Temperature (LST) represents the skin temperature of the land, expressed in Kelvin. Unlike other vegetation indices, LST is available in CLMS as an hourly variable. However, due to limitations in the temporal extents of individual datasets, three complementary sources were used to ensure full temporal coverage of the IberFire datacube.
For the period from 01/12/2007 to 10/06/2010, LST values were obtained from the aggregated ERA5-Land dataset, which provides global daily average skin temperature estimates at a spatial resolution of 9 km $\times$ 9 km. From 11/06/2010 to 18/01/2021, a CLMS source offering hourly LST at a resolution of 5km $\times$ 5km was used. For each timestamp, the daily average value was calculated from retrieved hourly values. Finally, from 19/01/2021 to 31/12/2024, a second CLMS source was used, also with hourly values with 5km $\times$ 5km resolution, and the same daily temporal averaging method was applied.
Finally, to harmonise all three datasets with the 1km $\times$ 1km resolution of IberFire, nearest-neighbour resampling was applied in each case.
@@
The Soil Water Index (SWI) is a moisture-related indicator that estimates the percentage of water retained in the upper layers of the soil. It is derived from observations of Surface Soil Moisture (SSM) using an exponential filtering approach, giving more weight to recent measurements while smoothing the temporal signal [29].
@@
To capture moisture dynamics at different temporal scales, four SWI variants were downloaded and included in the datacube, corresponding to $T=1,5,10,20$ and named SWI_001, SWI_005, SWI_010, SWI_020 respectively. The data were retrieved from CLMS, which provides daily SWI values at a spatial resolution of 12.5km $\times$ 12.5km. To align with the datacube’s 1 km $\times$ 1 km resolution, nearest-neighbour interpolation was applied during the resampling process.

### v2

@@ 2.5 Land usage: Corine Land Cover
The Copernicus Corine Land Cover (CLC) dataset [CLC] offers a standardised classification of land cover types across Europe, distinguishing 44 discrete categories. It represents the continent as a regular grid of 100m $\times$ 100m cells, assigning to each cell an integer value between 1 and 44 that corresponds to a specific land cover class.
@@ 44 Water bodies Marine waters Sea and ocean
The original spatial resolution of the CLC dataset is 100m $\times$ 100m; consequently, resampling of the data was needed to match the 1km $\times$ 1km resolution of the IberFire datacube. Using QGIS, the proportion of each of the 44 classes within each 1km $\times$ 1km cell was calculated. This resulted in 44 features, denoted as CLC_i for $i=1,2,\dots,44$, with values ranging between 0 and 1.
Given that these features represent proportions and therefore sum up to 1 for each 1km $\times$ 1km cell, and also considering the hierarchical structure of the CLC classification, five additional features were derived corresponding to the higher-level groupings defined in Label 1. Specifically, for each higher-level category, its proportion within a cell was computed by summing the relevant CLC_i variables that fall inside the category. For example, the proportion of Artificial surfaces was obtained by aggregating the values of CLC_1 through CLC_11.
@@
The CLC dataset is updated every six years; accordingly, the 2006 [CLC_2006_version], 2012 [CLC_2012_version], and 2018 [CLC_2018_version] editions were used in this work, as the 2024 version was not yet available at the time of writing. As previously discussed, time-invariant features in the datacube can be stored using only spatial dimensions. Given the low update frequency of the CLC dataset, including a temporal dimension would result in unnecessary duplication of values. Therefore, all CLC-derived features were stored without the time axis. To differentiate between editions, the corresponding year was appended to each feature name, as better described in Table 7 of Section 3. Each of the three CLC editions contributed 63 variables, this approach ultimately yielded $63\cdot 3=189$ distinct CLC-derived spatial-only features in the datacube. However, since only one CLC edition is relevant for any given $(x,y,t)$ cell, these 189 variables effectively represent just 63 unique features. The selection of the appropriate CLC edition for each cell is described in detail in Section 5.
@@ 2.6 Topography variables: European Digital Elevation Model
Topography is a well-established factor influencing the spread and intensity of forest fires. To capture these effects, topographic features were downloaded from the European Digital Elevation Model (EU-DEM) provided by OpenTopography [EU_DEM]. The original EU-DEM dataset offers elevation values at a $\times30m$ spatial resolution. Using this elevation data, additional variables such as slope, aspect and roughness of the terrain can be derived, all of which are also offered by OpenTopography.
However, the specific methods used by OpenTopography to calculate the slope and roughness are not documented, leaving the units of these variables undefined. In contrast, the aspect is expressed in degrees, ranging from 0°to 360°, indicating the direction in which the slope of each $\times30m$ terrain cell faces, with 0° corresponding to the north and the values increasing clockwise.
The elevation, slope, roughness and aspect were downloaded and, for the first three features, the mean and standard deviation in each $\times1km$ cell were calculated. Therefore, a total of 6 different features were obtained from elevation, slope, and roughness. For aspect, the feature was discretised into 8 different classes defined in Table 4. For each of the $\times1km$ cell, the proportion of each class inside the cell was calculated, resulting in 8 new features called aspect_i for $i=1,\dots,8$. Additionally, an extra feature, aspect_NODATA, was included to represent the proportion of pixels within each $\times1km$ cell for which no aspect value was available, which is highly correlated with is_waterbody. This lack of data typically occurs in areas covered by lakes and rivers, where aspect values are undefined due to the absence of terrain elevation gradients.
@@ 2.7 Human activity
It is widely recognised that most wildfires are directly related to human activities, either intentionally or accidentally [liz-lopez_spain_2023]. To model this phenomenon, six human-related explanatory variables were considered: distance to roads, distance to waterways, distance to railways, designation within the Natura 2000 protected network [miteco_rednatura2000], population density, and holiday periods. These variables were used to derive a total of 20 features, which were subsequently integrated into the datacube.
@@
The first two variables, distance to roads [distance_to_roads] and distance to waterways [distance_to_waterways], were obtained from WorldPop [worldpop], which provides data at an approximate spatial resolution of $\times100m$, represented in kilometres. To upscale the data to the target resolution, the mean and standard deviation within each $\times1km$ cell were computed using QGIS. These aggregated statistics were then incorporated into the datacube, resulting in four derived features from the original two variables. These features were considered invariant over time and therefore added with only spatial coordinates.
@@
For distance to railways, railway vectorial data in Spain were retrieved from OpenStreetMap [openstreetmap]. Using QGIS, a $\times100m$ raster layer was generated to represent the distance to the nearest railway geometry from each $\times100m$ cell. Then, the same procedure applied to the previous two variables was then applied here as well, and the average and standard deviation were calculated. These two features were then added to the datacube as spatial data.
@@
The Natura 2000 network, a European ecological network for biodiversity conservation, was included due to its relevance to fire data reported by EFFIS. In the vector fire data provided by EFFIS, the proportion of each forest fire that occurred within Natura 2000 protected areas is specified. Using vectorial boundaries of the Natura 2000 network, a binary raster layer was created with QGIS indicating whether each $\times1km$ cell falls within the protected area. Since this variable remains constant over time, it was integrated as a spatial layer into the datacube.
@@
Population density was incorporated as a proxy for human presence and potential ignition sources. Annual data were obtained from WorldPop for the years 2008 to 2020 [popdens], due to the unavailability of more recent data. The original data were retrieved with a resolution of approximately 1km $\times$ 1km, and average resampling was applied to match the coordinate values of IberFire. Given the annual update frequency of the data, population density were stored as spatial features to prevent unnecessary duplication of values, mirroring the process conducted with the CLC-derived features. Consequently, this variable appears as 13 spatial layers in the dataset, but functionally represents a single feature, as only the population density for the relevant year should be selected for each $(x,y,t)$ cell.
@@
Finally, a binary feature representing holiday periods was included. This variable accounts for the increased presence of people in rural and natural areas during weekends and public holidays, potentially raising the risk of fire ignition. A cell was flagged as a holiday if the corresponding date was a Saturday, Sunday, or a national or regional public holiday in Spain. To determine public holidays, the Python library holidays [arkadii_yakovets_2025_15169945] was utilised. This library provides holiday dates for various countries and their subdivisions, allowing for precise identification of holidays at the autonomous community level in Spain. Consequently, the AutonomousCommunities feature was employed to assign the appropriate regional holidays to each cell. The resulting is_holiday binary layer was added as a spatio-temporal feature within the datacube.
@@ 2.8 Meteorological variables: ERA5-Land
To account for these dynamics, data from ERA5-Land [ERA5_validation] were integrated, a high-resolution global reanalysis dataset provided by the Copernicus Climate Data Store. ERA5-Land offers hourly meteorological variables at a spatial resolution of 9 km $\times$ 9 km, from 1950 to the present. These variables are obtained as a combination of meteorological measurements and numerical weather prediction models through data assimilation techniques. This approach ensures spatial and temporal consistency.
ERA5-Land is particularly well-suited for environmental modelling due to its global coverage, temporal consistency, and availability of multiple curated atmospheric variables relevant to fire risk assessment. Within the Copernicus Climate Data Store, two data access modes are available: raw hourly observations [era5_hourly_data] and post-processed daily statistics [era5_postprocessed_data]. Both sources were used in the construction of the IberFire datacube, depending on the specific requirements of each variable.
Although ERA5-Land provides temporally consistent data, its availability is subject to a delay of five days, making it unsuitable for real-time fire risk prediction, where daily forecasts are required. Consequently, while ERA5-Land data were used to construct the datacube, real-time model deployment is intended to rely on meteorological station measurements.
The use of meteorological station measurements for the construction of the datacube was also considered, but the lack of historical data for many variables in various meteorological stations posed significant limitations.
@@
A total of 17 meteorological features were derived from ERA5-Land data. The features can be grouped as follows: temperature, relative humidity, surface pressure, precipitations, wind speed, and wind direction. This subsection outlines the methodology used to obtain and incorporate these variables into the IberFire datacube, including the selection of source variables and the transformations necessary to ensure compatibility with the format and units used in AEMET observations.
@@
Relative humidity is the percentage of moisture in the air relative to the maximum amount the air can hold at a given temperature. It is a critical variable for fire risk assessment, as it directly affects the moisture content of vegetation and, consequently, the likelihood of ignition and propagation. However, ERA5-Land does not provide this variable directly, while AEMET does. To address this, hourly relative humidity values were derived from two available ERA5-Land variables: the 2m temperature and the 2m dewpoint temperature. The last one represents the temperature at which the air becomes saturated with moisture, at 2 metres above the ground. Using these two temperature variables, relative humidity values were computed by applying the Magnus formula [ImprovedMagnusFormApproximationofSaturationVaporPressure], an empirically validated approach for estimating saturation vapor pressure:
$\textit{Relative Humidity}=\frac{\exp(\frac{17.625\cdot D_{p}}{243.04+D_{p}})}{\exp(\frac{17.625\cdot T}{243.04+T})}$ (1)
@@
From the computed hourly values, four daily statistics were derived and included in the datacube: the mean, minimum, maximum, and range of relative humidity.
Computing daily relative humidity statistics first required deriving hourly relative humidity values from the corresponding hourly temperature and dewpoint temperature data. This step was essential because daily statistics of relative humidity cannot be accurately obtained from the summary statistics of the temperature variables provided by ERA5-Land. For instance, inserting the daily mean values of $D_{p}$ and $T$ into Equation (1) does not yield the correct daily mean of relative humidity.
@@
$\lVert\vec{(u,v)}\rVert=\sqrt{\lvert{u}\rvert^{2}+\lvert{v}\rvert^{2}}\,.$ (2)
@@
wind_speed_mean. The data were stored with the original ERA5-Land units, metres per second (m/s), since they match with the units used by AEMET.
@@
Wind direction is conventionally defined as the direction from which the wind originates, expressed in degrees measured clockwise from true north. AEMET provides two daily wind-direction metrics: the average wind direction and the direction at the daily maximum wind speed. Meanwhile, ERA5-Land offers hourly wind data at a resolution of 9km $\times$ 9km, represented by horizontal (u_wind) and vertical (v_wind) components. To convert these components into standard wind direction, equation (3) is employed, where $\alpha$ denotes the angle between $\vec{(u,v)}$ and the north vector $\vec{(0,1)}$.
$\textit{Wind direction}=(\alpha-180^{\circ})\bmod 360^{\circ}\,,\quad\alpha=\angle\bigl(\vec{(0,1)},\vec{(u,v)}\bigr)\,.$ (3)
@@ 2.9 Vegetation indices: Copernicus Land Monitoring Service
The last group of features integrated into the datacube is the vegetation indices, which are highly used for making fire-risk predictions, as they can represent the dryness of the plant life. Five different indices were added: FAPAR (Fraction of Photosynthetically Active Radiation), LAI (Leaf Area Index), NDVI (Normalised Difference Water Index), LST (Land Surface Index), and SWI (Soil Water Index). All of them were retrieved from the Copernicus Land Monitoring Service (CLMS) [CopernicusLandMonitoringService] at various spatial and temporal resolutions.
For each vegetation index, multiple data sources from the Copernicus Land Monitoring Service (CLMS) were employed, each differing in spatial and temporal resolution. Depending on the source, some datasets provide global coverage, while others are limited to the European continent. To ensure efficiency and relevance, the data were downloaded specifically for the region of interest illustrated in Figure 1.
In the following, a detailed description is provided of the sources selected for each index, the procedures followed during data acquisition, and the preprocessing steps applied to harmonise all variables with the target spatial resolution of 1km × 1km and the daily temporal granularity required by the IberFire datacube. This process was done using Python, and all the vegetation indices were incorporated into the dataset as spatio-temporal features.
@@
The Fraction of Absorbed Photosynthetically Active Radiation (FAPAR) is a key biophysical variable that is commonly used as an indicator of vegetation health. It has been identified as one of the 50 Essential Climate Variables by the Global Climate Observing System (GCOS) [GCOS]. In the context of wildfire risk prediction, FAPAR is particularly relevant as it reflects the photosynthetic activity and water stress level of vegetation, which are tightly linked to fuel dryness.
The CLMS provides multiple datasets for the FAPAR variable; however, no single source spans the entire temporal range required by the IberFire datacube. To achieve full temporal coverage, two complementary datasets were integrated. The first source [FAPAR_v1] offers FAPAR measurements at a spatial resolution of 1km $\times$ 1km with a 10-day temporal frequency and was used to cover the period from 01/12/2007 to 30/04/2020. The second source [FAPAR_v2], used for the remaining period from 01/05/2020 to 31/12/2024, provides enhanced spatial resolution at 300m $\times$ 300m, while maintaining the same temporal frequency.
The two datasets were harmonised as follows. For the first source, since its native resolution matched the target resolution of the datacube, values were directly aligned to the grid using nearest-neighbour resampling. In contrast, the higher-resolution values from the second source were aggregated by computing the mean FAPAR value within each corresponding 1km $\times$ 1km cell, effectively applying the average resampling method.
@@
The download and preprocessing procedures for this variable followed the same approach as for the FAPAR index. A dataset with a 1 km spatial resolution was used for the period from 01/12/2007 to 30/04/2020 [LAI_v1] and was aligned to the datacube using nearest-neighbour resampling. For the remaining period, from 01/05/2020 to 31/12/2024, a dataset at 300 m resolution was employed [LAI_v2], and values were aggregated to 1 km resolution using average resampling.
@@
The Normalised Difference Vegetation Index (NDVI) is a remote sensing indicator that estimates vegetation health. It is calculated from the red and near-infrared spectral bands and typically ranges from 1-1 to $1$, with higher values indicating denser and healthier vegetation. In the context of wildfire risk assessment, NDVI serves as a proxy for vegetation condition and fuel availability.
As with the previous indices, two complementary datasets from the CLMS were used to ensure full temporal coverage. The first dataset [NDVI_v1], with a spatial resolution of $1km\times 1km$ and a 10-day frequency, covers the period from 01/12/2007 to 30/06/2020. The second dataset [NDVI_v2], with a finer resolution of 300 m and the same temporal frequency, was used from 01/07/2020 to 31/12/2024. The harmonization process mirrored that used for FAPAR and LAI: nearest-neighbour resampling was applied to the 1 km dataset, while the 300 m data were aggregated to 1 km using average resampling.
@@
Land Surface Temperature (LST) represents the skin temperature of the Earth’s surface, expressed in Kelvin. Unlike other vegetation-related indices, LST is available from CLMS as an hourly variable. However, to achieve full temporal coverage for the 2007–2024 period in the IberFire datacube, three complementary data sources were integrated, as no single dataset spans the entire time range. Specifically, the ERA5-Land skin temperature variable was used to fill in the earlier years, during which CLMS does not provide LST data.
For the period from 01/12/2007 to 10/06/2010, LST values were obtained from the aggregated ERA5-Land dataset [era5_postprocessed_data], which provides global daily average skin temperature values at a spatial resolution of 9 km $\times$ 9 km. From 11/06/2010 to 18/01/2021, a CLMS source offering hourly LST at a resolution of 5km $\times$ 5km was used [LST_v2]. For each daily timestamp, the average value was calculated from retrieved hourly values. Finally, from 19/01/2021 to 31/12/2024, a second CLMS source was used [LST_v3], also with hourly values with 5km $\times$ 5km resolution, and the same daily temporal averaging method was applied.
Finally, to harmonise all three datasets with the 1km $\times$ 1km resolution of IberFire, nearest-neighbour resampling was applied in each case, since the target resolution is finer than the original resolutions.
@@
The Soil Water Index (SWI) is a moisture-related indicator that estimates the percentage of water retained in the upper layers of the soil. It is derived from observations of Surface Soil Moisture (SSM) using an exponential filtering approach, giving more weight to recent measurements while smoothing the temporal signal [SWI_formula].
@@
To capture moisture dynamics at different temporal scales, four SWI variants were downloaded and included in the datacube, corresponding to $T=1,5,10,20$ and named SWI_001, SWI_005, SWI_010, SWI_020 respectively. The data were retrieved from CLMS, which provides daily SWI values at a spatial resolution of 12.5km $\times$ 12.5km [SWI_dataset]. To align with the datacube’s 1 km $\times$ 1 km resolution, nearest-neighbour interpolation was applied during the resampling process.